# From Frege to chatGPT: Compositionality in language, cognition, and deep neural networks

### v1

@@ 4.4 In-Context Compositionality via Pretraining
The original tasks used to evaluate compositionality (e.g. the SCAN dataset; Lake and Baroni,, 2018) were developed at a time when most models were trained from scratch to perform a specific task, such as machine translation (Bahdanau et al.,, 2014; Vaswani et al.,, 2017). However, researchers soon began to utilize self-supervised methods that did not require human labels (e.g., annotations, translations, etc.), thus allowing models to be trained on the huge amounts of unlabeled text available on the internet. The dominant self-supervised task has been language modeling, where a neural network parameterizes a distribution over sequences of natural language text. For example, in causal language modeling this distribution is expressed in terms of the probability of the next token given the entire preceding context. Researchers found that as transformer models were scaled up (in terms of number of parameters and amount of data) language modeling performance continued to improve. This meant that better performance could be achieved if large language models were pretrained on unlabeled text and then fine-tuned on the specific tasks of interest, rather than training from scratch on the specific task (Devlin et al.,, 2019; Peters et al.,, 2018; Radford et al.,, 2019).
@@
Researchers then discovered that at even larger scales (e.g., at the scale of GPT-3, 175 billion parameters), you could in some cases forgo the fine-tuning step and simply ask these large language models to do a new task in context by providing a few examples of the task to be performed (Brown et al.,, 2020). For example, rather than fine-tuning a pretrained model to predict whether a human-labeled restaurant review was positive or negative, you could prompt the model with a few study examples consisting of reviews and their corresponding labels, and test it on a test query—an unlabeled review (also included in the same context at the end of the prompt). This ability to flexibly solve new tasks that were explicitly instructed or demonstrated with a few examples, termed in-context learning (ICL), was further improved when models were fine-tuned on human-generated instruction-following datasets (Wei et al., 2022a, ) and with reinforcement learning by human feedback (RLHF; Ouyang et al.,, 2022). This created yet another paradigm shift in natural language processing: very large models are trained on very large datasets and then tested in context on specific tasks.
@@
From a theoretical perspective, this “pretrain-test” paradigm (Hupkes et al.,, 2023) has important similarities with the metalearning setting described above. In metalearning, models acquire inductive biases by training on a distribution of tasks. The pretraining step in LLMs, where they are trained to predict the next word on a very large dataset, can be seen as training on a distribution of tasks (see Figure 8; Brown et al.,, 2020; Chan et al.,, 2022; Radford et al.,, 2019). Sometimes this prediction will look more like guessing based on frequencies (e.g., memorizing a common idiom), but other times the task can involve more sophisticated behaviors such as understanding recursion in computer programming (see Figure 8). There are doubtless many instances in these large datasets where the best way to predict the next word is to learn something in context, and perhaps even to compose novel concepts in context (see “in-context composition” in Figure 8). Thus, although it is easy to dismiss next-word prediction as a simplistic or purely statistical training objective, upon further reflection one can see how it might put pressure on the model to develop an inner-loop or ICL algorithm in its forward activation dynamics, and perhaps even to endow this inner-loop algorithm with the property of compositionality.
@@
LLMs at this scale perform quite well on many kinds of tasks that seem to require compositionality. As anyone who has interacted with chatGPT or similar products knows, these models are capable of writing whole paragraphs of coherent text about almost any topic, including novel concepts or made-up words that are included in the prompt. They can construct sentences with sophisticated grammatical structure (Linzen and Baroni,, 2021), and can put together well-reasoned arguments (Herbold et al.,, 2023). They even recapitulate some of the syntactic phenomena thought to require innate grammatical constraints such as syntactic island effects (Wilcox et al.,, 2023, although see Lan et al., 2024). They are also surprisingly adept at computer programming, and are capable of writing novel functions or scripts (e.g., involving compositions of known functions) to accomplish a goal given in natural language instructions. These models have achieved state-of-the-art performance on natural language inference tasks such as recognizing logical entailment (Brown et al.,, 2020), and have even demonstrated human-like tendencies in analogical reasoning tasks such as Raven’s Progressive Matrices (Webb et al.,, 2023), which arguably require compositionality as well.
@@
Researchers have also evaluated these models on the tasks designed specifically to test compositional generalization capabilities (Lake and Baroni,, 2018). As in the metalearning setting, the evaluation works slightly differently than in the original SCAN dataset, where models were trained from scratch on tens of thousands of examples without “jump” and tested on instructions containing “jump.” In the pretrain-test setting, the pretrained LLMs are given just a few study examples from the SCAN task in context and asked to generalize to new test samples that are also given in the same input. The compositional generalization test can be replicated by excluding certain kinds of examples from the context (e.g., longer examples, or more complicated constructions involving “jump”). When GPT-3 (code-davinci-002) was evaluated with a standard prompt on a compositional split of the SCAN task in this way, it only achieved 16.7% accuracy (Zhou et al., 2023a, ). [^21] Other studies have tested different versions of GPT with slightly different setups (e.g., slightly different prompts) and found similar results, although GPT-4 achieved 58% when the study examples given in the prompt were sorted by length (Lake and Baroni,, 2023). However, the inclusion of novel prompting methods improved performance to a near-perfect accuracy of 99.7%. These prompting methods, called “chain-of-thought” prompting (Wei et al.,, 2023) and “least-to-most” prompting (Zhou et al., 2023a, ), have been shown to improve performance on many reasoning tasks. In chain-of-thought prompting (Kojima et al.,, 2023; Wei et al.,, 2023), the model is asked to “think step-by-step,” and/or given examples that show intermediate steps of reasoning. In least-to-most prompting (Drozdov et al.,, 2022; Zhou et al., 2023a, ), these in-context examples are simply given in order of least difficult to most difficult. Many other prompting strategies have been developed to further improve performance on similar tasks (Huang and Chang,, 2023; Nye et al.,, 2021; Press et al.,, 2023; Yao et al.,, 2023).
@@
Figure 8: Predicting the next word on a large corpus of text can be seen as a kind of metalearning (Brown et al.,, 2020; Radford et al.,, 2019), in which the LLM trains on tasks requiring diverse capabilities such as memorization of factual knowledge, idioms, inflectional morphology (past tense), programming, and identifying the indirect object of a sentence (Wang et al.,, 2022). In some cases, successful prediction of the next word may require in-context composition of novel elements. Analogous to the metalearning scenario, a distinction can be made between the outer-loop learning algorithm training the weights of the network, and an implicit inner-loop learning algorithm that operates on information given in context (green) to make inferences relevant to predicting the next word(s) (blue).
Taken together, the impressive performance of LLMs shows that at least in some cases, compositionality can arise as an emergent property of the ICL algorithm implemented in the forward activation dynamics of a large model trained to predict the next word on a large corpus of text (and in some cases also instruction-tuned and trained with reinforcement learning by human feedback). Similar to the studies using metalearning models (Lake,, 2019; Lake and Baroni,, 2023), these LLMs demonstrate that even though the outer-loop learning algorithm may not itself be capable of compositional (o.o.d.) generalization, it can endow an inner-loop, ICL algorithm with this capability. This again shows that the o.o.d. generalization problem solved by the inner-loop ICL algorithm may be an i.i.d. generalization problem from the perspective of the outer-loop learning algorithm, which has probably had many opportunities to learn how to learn similar problems in context from its training set (see Figure 8).
Can these findings about LLMs inform us about human compositionality? It is of course difficult to make any inferences from LLMs to humans because there are so many differences between them. For example, transformers are not biologically plausible in many ways, the scale of the training dataset is orders of magnitude larger than a human lifetime’s worth of language data, etc. (Frank,, 2023; Linzen and Baroni,, 2021; Pavlick,, 2023; Warstadt and Bowman,, 2024). However, these results do show that in principle it is possible to metalearn an ICL algorithm capable of compositional generalization by simple self-supervised learning in the outer-loop. Furthermore, as opposed to the experiments with explicit metalearning (Lake and Baroni,, 2023), the impressive compositional behaviors of LLMs show that the distribution of tasks on which a metalearner is trained doesn’t need to be deliberately contrived to encourage compositionality—strong compositional generalization abilities emerge in these models even when they are trained on unstructured text from the internet. This suggests that humans, who also learn from relatively unstructured experiences, could plausibly metalearn throughout development to implement a compositional inner-loop learning algorithm as well (Russin et al., 2024a, ).
@@ 5 Discussion
Even as deep neural networks blew past competing frameworks in virtually every area of AI, initial investigations into their compositionality seemed to confirm traditional intuitions that they were not capable of replicating the kinds of compositional generalization that humans exhibit (see section 4.1; Lake and Baroni,, 2018; Kim and Linzen,, 2020). One way of understanding these initial findings is from the statistical learning perspective: compositional generalization requires out-of-distribution (o.o.d.) generalization or extrapolation, but neural networks assume that train and test data will be independent and identically distributed (i.i.d.). Standard neural networks fail on these problems, then, because they lack the strong inductive biases that would allow them to generalize outside of their training distributions. Some have taken these results to demonstrate that neural networks will never achieve compositionality on their own and concluded that they must therefore be augmented with classical symbolic processes (“neuro-symbolic hybrids,” see section 4.2; Marcus and Davis,, 2020; Marcus,, 2020).
More recent advances complicate this picture, suggesting alternative ways in which strictly neural models might explain the same generalization behaviors (Baroni,, 2020). These advances include (1) architectural inductive biases (section 4.2), where a specialized architecture endows the network with an inductive bias for compositionality, (2) metalearning (section 4.3), where a network is trained on many compositional generalization problems so that it learns to implement a compositional inner-loop learning algorithm, and (3) large pretrained language models (section 4.4), which also implement an emergent compositional inner-loop learning algorithm, but come to do so via large-scale pretraining on next-word prediction.
These methods for imparting inductive biases for compositionality into neural networks overcome the challenge of o.o.d. generalization in different ways. Architectural inductive biases that separate structure-related processing from content-related processing (Russin et al., 2020a, ; Behrens et al.,, 2018; Schlag et al.,, 2019; Webb et al.,, 2024) can be seen as transforming an o.o.d. compositional generalization problem into two i.i.d. generalization problems—one involving structure and the other involving content (see Figure 4; Russin et al., 2020a, ). Factorizing structure from content in this way allows generalization to any novel structure-content combination, as long as the particular structures are familiar. The metalearning and large pretraining settings offer a different perspective on the challenge of o.o.d. generalization. In both of these cases, what looks like an o.o.d. problem for an inner-loop or in-context learning algorithm can be seen as i.i.d. for the outer-loop learning algorithm that trained the network in the first place (see Figures 6 and 7); Lake and Baroni,, 2023).
@@
More generally, these findings provide novel perspectives on human compositionality, suggesting that architectural inductive biases or metalearning may be important for explaining key aspects of these generalization behaviors in humans. In the following, we briefly discuss how these two perspectives can provide novel empirical predictions about the neural mechanisms and development of human compositionality.
5.1 Implications for Neural Mechanisms
Architectural inductive biases The strictly neural approach to architectural inductive biases assumes that there are important innate (non-classical) organizational features of the architecture of the human brain that encourage compositional generalization (Frankland and Greene,, 2020; O’Reilly et al.,, 2014; Russin et al., 2020b, ). Much of the work investigating such architectural inductive biases has explored principles related to the separation of structure and content (Smolensky et al., 2022a, ; Russin et al., 2020a, ; Frankland et al.,, 2021; Webb et al.,, 2024; Whittington et al.,, 2020).
It has been hypothesized that a similar separation constitutes a key organizational principle of the brain (Behrens et al.,, 2018; O’Reilly et al.,, 2021; Summerfield et al.,, 2020). In the visual system, the ventral visual stream (the “what” pathway) encodes object-related information and the dorsal visual stream (the “where” pathway) encodes spatial information (Ungerleider and Mishkin,, 1982). This organization, which has been sculpted over the course of evolutionary history (Mishkin and Ungerleider,, 1982), may encourage generalization to new structure-content (i.e., location-object) combinations in the visual domain (Pylyshyn,, 1989).
More generally, dorsal areas in the parietal cortex seem to encode structure-related information (Summerfield et al.,, 2020), including spatial (Black and Strub,, 1976; Chechlacz and Humphreys,, 2014; Friedman-Hill et al.,, 1995), temporal (Jazayeri and Shadlen,, 2015), and relational (Chafee,, 2013; Chafee et al.,, 2007) information, while ventral areas in the temporal lobe encode content-related information, including information about objects (Reddy and Kanwisher,, 2006) and semantic features (Visser et al.,, 2010). This gross separation between structure and content may extend into memory-related cortical networks (O’Reilly et al.,, 2021; Ranganath and Ritchey,, 2012) as well as the entorhinal cortex and hippocampus (Behrens et al.,, 2018; Manns and Eichenbaum,, 2006; Park et al.,, 2020; Russin et al., 2021b, ; Whittington et al.,, 2020; Zolfaghar et al.,, 2022), where structure-related information represented in medial entorhinal cortex and content-related information represented in lateral entorhinal cortex are bound in hippocampus (Behrens et al.,, 2018; Manns and Eichenbaum,, 2006). This organization of memory-related areas may facilitate generalization to novel structure-content combinations in memory-related domains (Whittington et al.,, 2020).
In addition to this broad separation between structure and content, there are also thought to be specific mechanisms for encouraging abstraction and content-independent processing in the prefrontal cortex (Badre et al.,, 2010; Frank and Badre,, 2012; Kriete et al.,, 2013; O’Reilly and Frank,, 2006; Rougier et al.,, 2005; Russin et al., 2020b, ; Wallis et al.,, 2001). For example, the prefrontal cortex has specialized gating mechanisms known to be important for working memory (O’Reilly and Frank,, 2006), and which may be involved in the uniquely human capacity for rule-learning and indirection (Rougier et al.,, 2005; Kriete et al.,, 2013). Furthermore, the prefrontal cortex is thought to learn sequential, top-down attentional procedures for accomplishing specific goals (Miller and Cohen,, 2001; O’Reilly et al.,, 2020). If this learning is isolated from the particular objects or contents that are being attended (similar to the isolation in Figure 4), this may encourage content-independent representations and compositional generalization to emerge.
Unlike state-of-the-art neural networks like the transformers used in LLMs, the brain is a highly heterogeneous neural network architecture with many specialized features. Architectural motifs like the broad separation between structure-related and content-related processing and specialized circuits in prefrontal cortex seem like promising candidates for helping to explain human compositionality, but much work remains to be done to establish the precise contributions made by each of these proposed mechanisms to human generalization behaviors.
@@ 5.1 Implications for Neural Mechanisms
Metalearning The metalearning approach hypothesizes that human compositionality is a property of an inner-loop, in-context learning algorithm that has been metalearned by an outer-loop, in-weight learning algorithm (Binz et al.,, 2023; Dubey et al.,, 2020; Lake and Baroni,, 2023; Russin et al., 2024a, ; Wang et al.,, 2018). [^22] Some have assumed that metalearning in neural networks is a better model of learning on an evolutionary timescale than within an individual’s lifetime (Binz et al.,, 2023). Here, we assume that metalearning can take place in the brain over the course of an individual’s lifetime, and that in-context learning and in-weight learning can simultaneously exist within a single system. This idea is directly investigated by Russin et al., 2024b . This hypothesis predicts that when the neural mechanisms supporting in-context learning, rather than in-weight learning, are driving behavior, this should facilitate compositional generalization.
@@
Russin et al., 2024b found that this interplay between in-context and in-weight learning in neural network models explained human compositional generalization performance on a task recently studied in humans (Dekker et al.,, 2022). Both LLMs and metalearning neural networks achieved much better compositional generalization performance on the task when in-context learning rather than in-weight learning was responsible for behavior. Furthermore, Russin et al., 2024b found that the interplay between in-context and in-weight learning (Chan et al.,, 2022), which has similarities to a tradeoff between working memory and reinforcement learning observed in humans (Collins and Frank,, 2018; Rac-Lubashevsky et al.,, 2023), also explained the curriculum effects observed in the human study (Dekker et al.,, 2022).
@@
There is further support from cognitive and computational neuroscience for the idea that in-context learning and compositional generalization share a set of underlying neural mechanisms in humans. The prefrontal cortex is important for basic capacities known to be involved in human in-context learning, such as working memory (O’Reilly and Frank,, 2006), top-down attention (Cohen et al.,, 1990), and cognitive control (Miller and Cohen,, 2001), and, as previously mentioned, is thought to be important for compositional generalization processes such as rule learning (Calderon et al.,, 2022; Collins and Frank,, 2013; Frank and Badre,, 2012; Kriete et al.,, 2013), reasoning (Crescentini et al.,, 2011; Goel,, 2007), and processing complex syntax (Thompson-Schill,, 2005). Furthermore, it has been hypothesized that metalearning (specifically, meta-reinforcement learning) is an important aspect of the functioning of the prefrontal cortex (Hattori et al.,, 2023; Wang et al.,, 2018).
@@
Thus, while it may be difficult to directly establish that human compositional generalization behaviors are emergent consequences of an inner-loop, in-context learning algorithm, there is indirect empirical support for the idea that compositional generalization and in-context learning share an underlying set of neural mechanisms in humans.
5.2 Implications for Human Development
Architectural inductive biases When taken seriously as models of human learning, the architectural inductive biases reviewed in section 4.2 generate concrete proposals about the emergence of compositionality over the course of human development. Both neuro-symbolic hybrids (Marcus,, 2020) and strictly neural architectural inductive biases (O’Reilly et al.,, 2014; Russin et al., 2020a, ; Webb et al.,, 2024) assume that humans are not perfectly unbiased learners (blank slates), and that innate constraints will ultimately play an important role in explaining the kinds of generalization behaviors they exhibit. The key difference between the two is that the former assumes innate representations that are guaranteed to compose, while the latter builds in innate architectural motifs, and then relies on learning to discover the most useful representations to accomplish a given objective, allowing compositionality to emerge over the course of training (Bates et al.,, 1996). The strength of the neuro-symbolic hybrid approach (and of any approach to learning that assumes symbolic or compositioanl representational elements; Goodman et al.,, 2015; Piantadosi,, 2021; Quilty-Dunn et al.,, 2023; Yang and Piantadosi,, 2022) is that it provides a better guarantee of compositionality, but the strength of the strictly neural approach is that it allows greater flexibility to learn from data (Donatelli and Koller,, 2023).
The neuro-symbolic hybrid approach predicts that the generalizations humans make are universally constrained by innate representations, such as grammatical principles (Chomsky,, 1965; Lightfoot,, 1989; Montague,, 1970; Pinker,, 1994; Pinker and Bloom,, 1990) or the available logical primitives (Piantadosi et al.,, 2016). The strictly neural approach, on the other hand, predicts that these generalizations will be determined to a large extent by prior learning experiences (Prince and Smolensky,, 2004). For example, if the brain embodies an architectural inductive bias to separate structure-related processing from content-related processing, then it should be more difficult to generalize to completely novel structures than it is to generalize to novel compositions of familiar structures/contents. There has been a great deal of work investigating the degree to which innate constraints or core knowledge contribute to cognitive development (Carey,, 2011; Spelke and Kinzler,, 2007; Spelke,, 2022; Feiman et al.,, 2022), but few studies have sought to directly test fine-grained empirical predictions made by the specific architectural inductive biases shown to improve compositional generalization in modern neural networks. Future work could take current models and investigate where their inductive biases allow them to generalize compositionally and where they fail, comparing these results against human data. Ultimately, both hybrid and strictly neural approaches accept that humans are not maximally unbiased learners, and have specific innate inductive biases that facilitate compositional generalization.

### v2

@@ 3.3 Large Language Models
The original tasks used to evaluate compositionality (e.g. the SCAN dataset; Lake and Baroni,, 2018) were developed at a time when most models were trained from scratch to perform a specific task, such as machine translation (Bahdanau et al.,, 2014; Vaswani et al.,, 2017). However, researchers soon began to utilize self-supervised methods that did not require human labels (e.g., annotations, translations, etc.), thus allowing models to be trained on the huge amounts of unlabeled text available on the internet. The dominant self-supervised task has been language modeling, where a neural network parameterizes a distribution over sequences of natural language text. For example, in causal language modeling this distribution is expressed in terms of the probability of the next token given the entire preceding context. Researchers found that as transformer models were scaled up (in terms of number of parameters and amount of data) language modeling performance continued to improve. This meant that better performance could be achieved if LLMs were pretrained on unlabeled text and then fine-tuned on the specific tasks of interest, rather than training from scratch on the specific task (Devlin et al.,, 2019; Peters et al.,, 2018; Radford et al.,, 2019).
@@
Researchers then discovered that at even larger scales (e.g., at the scale of GPT-3, 175 billion parameters), you could in some cases forgo the fine-tuning step and simply ask these LLMs to do a new task in context by providing a few examples of the task to be performed (Brown et al.,, 2020). For example, rather than fine-tuning a pretrained model to predict whether a human-labeled restaurant review was positive or negative, you could prompt the model with a few study examples consisting of reviews and their corresponding labels, and evaluate it on a test query—an unlabeled review. These emergent ICL abilities allow LLMs to flexibly solve new tasks that are explicitly instructed or demonstrated with a few examples, and is further improved when models are fine-tuned on human-generated instruction-following datasets (Wei et al., 2022a, ) and with reinforcement learning by human feedback (RLHF; Ouyang et al.,, 2022). This created yet another paradigm shift in natural language processing: very large models are trained on very large datasets and then tested in context on specific tasks.
@@
From a theoretical perspective, this “pretrain-test” paradigm (Hupkes et al.,, 2023) has important similarities with the metalearning setting described above. In metalearning, models acquire inductive biases by training on a distribution of tasks. The pretraining step in LLMs, where they are trained to predict the next word on a very large dataset, can be seen as training on a distribution of tasks (see Figure 2; Brown et al.,, 2020; Chan et al.,, 2022; Radford et al.,, 2019). Sometimes this prediction will look more like guessing based on frequencies (e.g., memorizing a common idiom), but other times the task can involve more sophisticated behaviors such as understanding recursion in computer programming. There are doubtless many instances in these large datasets where the best way to predict the next word is to learn something in context, and perhaps even to compose novel concepts in context (see “in-context composition” in Figure 2). Thus, although it is easy to dismiss next-word prediction as a simplistic or purely statistical training objective, upon further reflection one can see how it might put pressure on the model to develop an ICL algorithm in its forward activation dynamics, and perhaps even to endow this ICL algorithm with the property of compositionality.
@@
LLMs at this scale perform quite well on many kinds of tasks that seem to require compositionality. As anyone who has interacted with chatGPT or similar products knows, these models are capable of writing whole paragraphs of coherent text about almost any topic, including novel concepts or made-up words that are included in the prompt. They can construct sentences with sophisticated grammatical structure (Linzen and Baroni,, 2021), and can put together well-reasoned arguments (Herbold et al.,, 2023). They even recapitulate some of the syntactic phenomena thought to require innate grammatical constraints such as syntactic island effects (Wilcox et al.,, 2023, although see Lan et al., 2024). They are also adept at computer programming, and are capable of writing novel functions or scripts (e.g., involving compositions of known functions) to accomplish a goal given in natural language instructions. These models have achieved state-of-the-art performance on natural language inference tasks such as recognizing logical entailment (Brown et al.,, 2020), and have even demonstrated human-like tendencies in analogical reasoning tasks such as Raven’s Progressive Matrices (Webb et al.,, 2023), which arguably require compositionality as well.
@@
Researchers have also evaluated these models on the tasks designed specifically to test compositional generalization capabilities (Lake and Baroni,, 2018). As in the metalearning setting, the evaluation works slightly differently than in the original SCAN dataset, where models were trained from scratch on tens of thousands of examples without “jump” and tested on instructions containing “jump.” In the pretrain-test setting, the pretrained LLMs are given just a few study examples from the SCAN task in context and asked to generalize to new test samples that are also given in the same input. The compositional generalization test can be replicated by excluding certain kinds of examples from the context (e.g., longer examples, or more complicated constructions involving “jump”). When GPT-3 (code-davinci-002) was evaluated with a standard prompt on a compositional split of the SCAN task in this way, it only achieved 16.7% accuracy (Zhou et al., 2023a, ). However, the inclusion of novel prompting methods improved performance to a near-perfect accuracy of 99.7%. These prompting methods, called “chain-of-thought” prompting (Wei et al.,, 2023) and “least-to-most” prompting (Zhou et al., 2023a, ), have been shown to improve performance on many reasoning tasks. In chain-of-thought prompting (Kojima et al.,, 2023; Wei et al.,, 2023), the model is asked to “think step-by-step,” and/or given examples that show intermediate steps of reasoning. In least-to-most prompting (Drozdov et al.,, 2022; Zhou et al., 2023a, ), these in-context examples are simply given in order of least difficult to most difficult. Many other prompting strategies have been developed to further improve performance on similar tasks (Huang and Chang,, 2023; Nye et al.,, 2021; Press et al.,, 2023; Yao et al.,, 2023).
@@
Figure 2: Predicting the next word on a large corpus of text can be seen as a kind of metalearning (Brown et al.,, 2020; Radford et al.,, 2019), in which the LLM trains on tasks requiring diverse capabilities such as memorization of factual knowledge, idioms, inflectional morphology (past tense), programming, and identifying the indirect object of a sentence (Wang et al.,, 2022). In some cases, successful prediction of the next word may require in-context composition of novel elements. As in metalearning, a distinction can be made between the IWL algorithm training the weights of the network, and an implicit ICL algorithm that operates on information given in context to make inferences relevant to predicting the next word(s).
Taken together, the impressive performance of LLMs shows that at least in some cases, the behavioral signatures of compositionality can emerge from the ICL algorithm implemented in the forward activation dynamics of a large model trained to predict the next word on a large corpus of text. Similar to the studies using metalearning models (Lake,, 2019; Lake and Baroni,, 2023), these LLMs demonstrate that even though the IWL algorithm may not itself be capable of compositional (o.o.d.) generalization, it can endow an emergent ICL algorithm with this capability. This again shows that the o.o.d. generalization problem solved by the ICL algorithm may be an i.i.d. generalization problem from the perspective of the original IWL algorithm, which has probably had many opportunities to learn how to learn similar problems in context from its training set (see Figure 2).
Can these findings about LLMs inform us about human compositionality? It is of course difficult to make any inferences from LLMs to humans because there are so many differences between them. For example, transformers are not biologically plausible in many ways, and the scale of LLM training datasets is orders of magnitude larger than a human lifetime’s worth of language data (Frank,, 2023; Linzen and Baroni,, 2021; Pavlick,, 2023; Warstadt and Bowman,, 2024). However, these results do show that in principle it is possible to metalearn an ICL algorithm capable of compositional generalization by simple self-supervised learning on natural language. Furthermore, in contrast with explicit metalearning (Lake and Baroni,, 2023), the impressive compositional behaviors of LLMs show that the distribution of tasks on which a metalearner is trained does not need to be deliberately contrived to encourage compositionality—strong compositional generalization abilities emerge in these models even when they are trained on unstructured text. This suggests that humans, who also learn from relatively unstructured experiences, could plausibly metalearn throughout development to implement a compositional ICL algorithm as well (Russin et al., 2024a, ).
@@ 4 Discussion
Even as deep neural networks blew past competing frameworks in virtually every area of AI, initial investigations into their compositionality seemed to confirm traditional intuitions that they were not capable of replicating the kinds of compositional generalization that humans exhibit (see section 3.1; Lake and Baroni,, 2018; Kim and Linzen,, 2020). One way of understanding these initial findings is from the statistical learning perspective: compositional generalization requires out-of-distribution (o.o.d.) generalization or extrapolation, but neural networks assume that train and test data will be independent and identically distributed (i.i.d.). Standard neural networks fail on these problems, then, because they lack the strong inductive biases that would allow them to generalize outside of their training distributions. Some have taken these results to demonstrate that neural networks will never achieve compositionality on their own and concluded that they must therefore be augmented with classical symbolic processes (“neuro-symbolic hybrids;” Marcus and Davis,, 2020; Marcus,, 2020).
More recent advances complicate this picture, suggesting alternative ways in which strictly neural models might explain the same generalization behaviors (Baroni,, 2020). In particular, recent results suggest that metalearning (section 3.2) and large-scale pretraining in LLMs (section 3.3) can endow neural network models with sophisticated in-context learning abilities, and that these abilities can in some cases capture compositional generalization behaviors.
@@
More generally, these findings provide a novel perspective on human compositionality, suggesting that metalearning and in-context learning may be important principles for explaining key aspects of these generalization behaviors in humans. In the following, we briefly discuss how these principles can provide novel empirical predictions about the neural mechanisms and development of human compositionality.
4.1 Metalearning and Neural Mechanisms
@@
The metalearning approach hypothesizes that human compositionality is a property of an in-context learning (ICL) algorithm that has been metalearned via the usual in-weight learning (IWL) (Binz et al.,, 2023; Dubey et al.,, 2020; Lake and Baroni,, 2023; Russin et al., 2024a, ; Wang et al.,, 2018). [^12] Some have assumed that metalearning in neural networks is a better model of learning on an evolutionary timescale than within an individual’s lifetime (Binz et al.,, 2023). Here, we assume that metalearning can take place in the brain over the course of an individual’s lifetime, and that ICL and IWL can simultaneously exist within a single system. This idea is directly investigated by Russin et al., 2024b . This hypothesis predicts that when the neural mechanisms supporting ICL, rather than IWL, are driving behavior, this should facilitate compositional generalization.
@@
Russin et al., 2024b found that this interplay between ICL and IWL in neural network models explained compositional generalization performance on a task recently studied in humans (Dekker et al.,, 2022). Both LLMs and metalearning neural networks achieved much better compositional generalization performance on the task when ICL rather than IWL was responsible for behavior. Furthermore, Russin et al., 2024b found that the dynamic interaction between ICL and IWL (Chan et al.,, 2022), which has similarities to a tradeoff between working memory and reinforcement learning observed in humans (Collins and Frank,, 2018; Rac-Lubashevsky et al.,, 2023), also explained the curriculum effects observed in the human study (Dekker et al.,, 2022).
@@
There is further support from cognitive and computational neuroscience for the idea that ICL and compositional generalization share a set of underlying neural mechanisms in humans. The prefrontal cortex is important for basic capacities known to be involved in human ICL, such as working memory (O’Reilly and Frank,, 2006), top-down attention (Cohen et al.,, 1990), and cognitive control (Miller and Cohen,, 2001), and, as previously mentioned, is thought to be important for compositional generalization processes such as rule learning (Calderon et al.,, 2022; Collins and Frank,, 2013; Frank and Badre,, 2012; Kriete et al.,, 2013), reasoning (Crescentini et al.,, 2011; Goel,, 2007), and processing complex syntax (Thompson-Schill,, 2005). Furthermore, it has been hypothesized that metalearning (specifically, meta-reinforcement learning) is an important aspect of the functioning of the prefrontal cortex (Hattori et al.,, 2023; Wang et al.,, 2018).
@@
Thus, while it may be difficult to directly establish that human compositional generalization behaviors are emergent properties of an ICL algorithm, there is indirect empirical support for the idea that compositional generalization and ICL share an underlying set of neural mechanisms in humans.
4.2 Metalearning and Human Development